# Geometry-Informed Neural Operator Transformer

### v4

@@ 3.2 Elasticity
For the elasticity example, we utilized the dataset generated by Li et al. [22] for training the Geo-FNO model. The dataset consists of unit cells of $[0,1]^{2}$ with arbitrary-shaped voids at the center, as illustrated in Fig. 3. The central void is defined by 45 points evenly distributed along the tangential direction, parameterized as
@@
Geo-FNO [22] employs 45 parametric void radii as geometry features, which limits its applicability to parametric geometries. In contrast, GINOT uses boundary points clouds to represent geometry, enabling it to handle arbitrary geometries. The stress solutions of the testing samples predicted by GINOT are shown in Fig. 3, highlighting the median and worst cases in terms of $L_{2}$ relative error. For the median case, the $L_{2}$ relative error is 1.23%, while the worst case has an $L_{2}$ relative error of 2.93%. The mean $L_{2}$ relative error across the testing samples is 1.34%, with a standard deviation of 0.46%.
@@ 3.3 Poisson equation
where $\Omega$ represents the domain, $u$ is the solution, $f$ is the source term, and $g$ is the Dirichlet boundary condition. For all samples, we set $f=1$ and $g=0$. Similar to the elasticity example, the domain $\Omega$ is parameterized using Eq. 7, with the radius $r_{i}\in[0.2,0.8]$ and $i=1,2,\ldots,144$. The geometries are illustrated in Fig. 4.
Two datasets were generated for this problem, each containing 6000 samples. The first dataset uses a structured mesh with a fixed 2737 nodes across all samples, while the second dataset employs an unstructured mesh, where the number of nodes varies between 1566 and 2838. The structured and unstructured meshes are illustrated in the last column of Fig. 4. For comparison with Geo-FNO, which does not support unstructured meshes with varying node counts, we focus on the structured mesh dataset. The Poisson equation was solved using FEM with FEniCSX. Of the total samples, 20% were reserved for testing, and the remaining 80% were used for training.
The comparison between GINOT predictions and the ground truth is presented in Fig. 4. For the unstructured mesh dataset, the median case has an $L_{2}$ relative error of 0.41%, while the worst case has an $L_{2}$ relative error of 2.27%. For the structured mesh dataset, only the median case is shown for brevity, with an $L_{2}$ relative error of 0.38%. The mean $L_{2}$ relative error for the testing samples of the structured mesh dataset is 0.45%, with a standard deviation of 0.17%, while for the unstructured mesh dataset, the mean $L_{2}$ relative error is 0.44%, with a standard deviation of 0.21%.
@@ 3.4 Bracket lug
We address the plasticity problem of bracket lugs using the dataset used in [24]. The bracket lugs, depicted in Fig. 5, are parameterized by three variables: length, thickness, and the radius of the circular hole. FEM simulations employ an elastic-plastic material model with linear isotropic hardening and quadratic hexahedral elements under small deformation assumptions. For training GINOT, the middle edge nodes of the quadratic hexahedral elements are removed. The surface mesh is extracted from the volume hexahedral elements, and its nodes are used as the boundary point cloud, which is then input to the geometry encoder of GINOT.
The dataset also includes variations in pressure loading across samples. To incorporate the loading parameter, we adopt the framework shown in Fig. 2. An MLP encodes the loading (shape of $B\times 1$) into a tensor with shape of $B\times d_{e}$, which is expanded to $B\times N_{s}\times d_{e}$ by repeating the tensor $N_{p}$ times, where $B$ is the batch size. This tensor is concatenated with the geometry encoder’s output and passed through an MLP, whose output serves as the KEY and VALUE for the attention mechanism. The dataset contains 3000 samples, with 20% reserved for testing.
The stress solutions for the testing samples predicted by GINOT are shown in Fig. 5, highlighting the median and worst cases based on $L_{2}$ relative error. The median case has an $L_{2}$ relative error of 0.41%, while the worst case reaches 3.78%. The mean $L_{2}$ relative error for the testing samples is 0.45%, with a standard deviation of 0.23%.
@@ 3.5 Micro-periodic unit cell
Figure 6: Mises stress and displacement solutions at the final strain step for the testing sample of the micro-periodic unit cell. The first row shows the median case ($L_{2}=10.37\%$), while the second row represents the 95th percentile case ($L_{2}=19.64\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column displays the true stress on the true deformed shape, the third column shows the predicted stress on the predicted deformed shape, and the last column highlights the absolute error of stress on the true deformed shape.
@@
The previous examples focus on 2D or 3D geometries with relatively simple shapes parameterized by a few variables. In this example, we consider micro-periodic unit cells (PUC) generated by the authors for metamaterial architecture design. These geometries are more complex and fully arbitrary. The dataset is created by first generating a 2D periodic Gaussian random field on a square domain $[0,1~{}\text{mm}]^{2}$. The Marching Squares algorithm is then applied to extract the contour at a random threshold value, defining the boundary of the 2D periodic unit cell, with field values below the threshold treated as void.
@@
The predicted solutions at the final strain step are shown in Fig. 6, comparing the ground truth with GINOT predictions. The first row highlights the median case with an $L_{2}$ relative error of 10.37%, while the second row shows the 95th percentile case with an $L_{2}$ relative error of 19.64%. The first column illustrates the input surface point cloud, the second column presents the true stress on the true deformed shape, the third column displays the predicted stress on the predicted deformed shape, and the last column shows the absolute error of stress on the true deformed shape. The mean $L_{2}$ relative error across the testing samples is 9.05%, with a standard deviation of 3.53%.
The stress and displacement solutions over historical strain steps are compared with the ground truth and illustrated in Fig. 7. For brevity, only the median case is presented here. Additional cases, such as the best and worst, can be found in our Github repository. The first row shows the true stress on the true deformed shape, the second row depicts the predicted stress on the predicted deformed shape, and the last row highlights the absolute error of stress on the true deformed shape. The overall $L_{2}$ relative error for historical steps is 9.08%. The mean $L_{2}$ relative error for the testing samples is 9.06%, with a standard deviation of 3.02%.
As detailed in Section 2.1, the geometry encoder in GINOT is designed to handle 1) nonuniform point cloud densities, 2) permutations in point order, and 3) padding within the point cloud. The first property is evident throughout the examples in this work. To demonstrate the latter two properties, we analyze the second-best test case in this example. Notably, GINOT maintains stable inference performance even when the point cloud is made sparser. In this analysis, the point cloud is shuffled, reduced to 80% of its original size, padded with 199 points at $(-1000,-1000)$, and shuffled again. These modified point clouds are input to the trained GINOT, with results shown in Fig. 8. The $L_{2}$ relative errors for the original, reduced and shuffled, padded, and shuffled padded point clouds are 4.24%, 4.65%, 4.24%, and 4.57%, respectively. The similar $L_{2}$ errors across these cases demonstrate the geometry encoder’s robustness to point order and padding. However, a substantial reduction in point cloud density can increase the $L_{2}$ error due to loss of geometric information, while shuffling and padding have negligible impact.
@@ 3.6 Jet engine bracket
The final example demonstrating GINOT’s capability is the Jet Engine Bracket (JEB) dataset [35]. This dataset comprises 2,138 samples of jet engine brackets derived from the "GE Jet Engine Bracket Challenge" guidelines. These geometries are complex 3D structures with arbitrary shapes and sizes. FEM simulations provide solution fields (e.g., stress, displacement) under four loading scenarios: vertical, horizontal, diagonal, and torsional, all with fixed loading magnitudes. In this example, we focus on the Mises stress solution under vertical loading.
@@
Figure 11: Mises stress solutions for the testing samples of the JEB dataset. The first row shows the best case ($\text{NRMSE}=0.74\%$, $L_{2}=16.9\%$), the second row represents the median case ($\text{NRMSE}=1.62\%$, $L_{2}=31.2\%$), and the third row depicts the worst case ($\text{NRMSE}=3.54\%$, $L_{2}=56.8\%$) based on NRMSE. The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
@@
The dataset includes CAD models, volume meshes in VTK format, and solution fields on fine quadratic tetrahedral elements. Instead of extracting surface nodes from the volume mesh, we extract the boundary point cloud directly from the CAD model vertices for each JEB sample using FreeCAD with PythonAPI, as illustrated in Fig. 9(a). To reduce computational costs, the volume mesh is downgraded from 10-node tetrahedrons to 4-node tetrahedrons by removing the middle nodes of each element’s edge. The nodes of these linear tetrahedrons are then used as query points in the solution decoder of GINOT, as shown in Fig. 9(a). Fig. 9(b) illustrates the number of boundary points in the cloud and the volume mesh nodes for the 2,138 JEB samples. The boundary point cloud contains between 2,253 and 10,388 points, while the volume mesh nodes range from 18,085 to 55,009.
@@
The dataset is divided into 90% for training and 10% for testing. Fig. 10 illustrates the training process of GINOT on the JEB dataset. The mean squared error (MSE) loss for both the training and validation datasets decreases steadily during training. However, the testing loss remains significantly higher than the training loss, suggesting overfitting. This behavior is expected given the complexity of the JEB dataset and its relatively small size.
@@
After training, we evaluate GINOT’s performance on both the training and testing samples. For the training samples, the mean $L_{2}$ relative error is 15.6% with a standard deviation of 3.6%, while for the testing samples, the mean $L_{2}$ relative error is 35.6% with a standard deviation of 9.88%. We also assess performance of the JEB dataset using the normalized root mean square error (NRMSE) defined in Eq. 5. The mean NRMSE for the training samples is 0.768% with a standard deviation of 0.167%, and for the testing samples, the mean NRMSE is 1.73% with a standard deviation of 0.525%. Fig. 11 illustrates the Mises stress solutions for the testing samples predicted by GINOT, showcasing the best, median, and worst cases based on NRMSE. The best case achieves an NRMSE of 0.74% and $L_{2}=16.9\%$, while the worst case has an NRMSE of 3.54% and $L_{2}=56.8\%$.

### v5

@@ 3.2 Elasticity
For the elasticity example, we utilized the dataset generated by li2023fourier for training the Geo-FNO model. The dataset consists of unit cells of $0,1]^{2}$ with arbitrary-shaped voids at the center, as illustrated in Fig.˜3. The central void is defined by 45 points evenly distributed along the tangential direction, parameterized as
@@
Geo-FNO [li2023fourier] employs 45 parametric void radii as geometry features, which limits its applicability to parametric geometries. In contrast, GINOT uses boundary points clouds to represent geometry, enabling it to handle arbitrary geometries. The stress solutions of the testing samples predicted by GINOT are shown in Fig.˜3, highlighting the median and worst cases in terms of $L_{2}$ relative error. For the median case, the $L_{2}$ relative error is 0.96%, while the worst case has an $L_{2}$ relative error of 2.31%. The mean $L_{2}$ relative error across the testing samples is 1.03%, with a standard deviation of 0.35%.
@@ 3.3 Poisson equation
where $\Omega$ represents the domain, $u$ is the solution, $f$ is the source term, and $g$ is the Dirichlet boundary condition. For all samples, we set $f=1$ and $g=0$. Similar to the elasticity example, the domain $\Omega$ is parameterized using Eq.˜6, with the radius $r_{i}\in[0.2,0.8]$ and $i=1,2,\ldots,144$. The geometries are illustrated in Fig.˜4.
Two datasets were generated for this problem, each containing 6000 samples. The first dataset uses a structured mesh with a fixed 2737 nodes across all samples, while the second dataset employs an unstructured mesh, where the number of nodes varies between 1566 and 2838. The structured and unstructured meshes are illustrated in the last column of Fig.˜4. For comparison with Geo-FNO, which does not support unstructured meshes with varying node counts, we focus on the structured mesh dataset. The Poisson equation was solved using FEM with FEniCSX. Of the total samples, 20% were reserved for testing, and the remaining 80% were used for training.
The comparison between GINOT predictions and the ground truth is presented in Fig.˜4. For the unstructured mesh dataset, the median case has an $L_{2}$ relative error of 0.41%, while the worst case has an $L_{2}$ relative error of 2.27%. For the structured mesh dataset, only the median case is shown for brevity, with an $L_{2}$ relative error of 0.38%. The mean $L_{2}$ relative error for the testing samples of the structured mesh dataset is 0.45%, with a standard deviation of 0.17%, while for the unstructured mesh dataset, the mean $L_{2}$ relative error is 0.44%, with a standard deviation of 0.21%.
@@ 3.4 Bracket lug
We address the plasticity problem of bracket lugs using the dataset used in [he2024geom]. The bracket lugs, depicted in Fig.˜5, are parameterized by three variables: length, thickness, and the radius of the circular hole. FEM simulations employ an elastic-plastic material model with linear isotropic hardening and quadratic hexahedral elements under small deformation assumptions. For training GINOT, the middle edge nodes of the quadratic hexahedral elements are removed. The surface mesh is extracted from the volume hexahedral elements, and its nodes are used as the boundary point cloud, which is then input to the geometry encoder of GINOT.
The dataset also includes variations in pressure loading across samples. To incorporate the loading parameter, we adopt the framework shown in Fig.˜2. An MLP encodes the loading (shape of $B\times 1$) into a tensor with shape of $B\times d_{e}$, which is expanded to $B\times N_{s}\times d_{e}$ by repeating the tensor $N_{p}$ times, where $B$ is the batch size. This tensor is concatenated with the geometry encoder’s output and passed through an MLP, whose output serves as the KEY and VALUE for the attention mechanism. The dataset contains 3000 samples, with 20% reserved for testing.
The stress solutions for the testing samples predicted by GINOT are shown in Fig.˜5, highlighting the median and worst cases based on $L_{2}$ relative error. The median case has an $L_{2}$ relative error of 0.41%, while the worst case reaches 3.78%. The mean $L_{2}$ relative error for the testing samples is 0.45%, with a standard deviation of 0.23%.
@@ 3.5 Micro-periodic unit cell
Figure 6: Mises stress and displacement solutions at the final strain step for the testing sample of the micro-periodic unit cell. The first row shows the median case ($L_{2}=7.54\%$), while the second row represents the 95th percentile case ($L_{2}=13.03\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column displays the true stress on the true deformed shape, the third column shows the predicted stress on the predicted deformed shape, and the last column highlights the absolute error of stress on the true deformed shape.
@@
The previous examples focus on 2D or 3D geometries with relatively simple shapes parameterized by a few variables. In this example, we consider micro-periodic unit cells (PUC) generated by the authors for metamaterial architecture design. These geometries are more complex and fully arbitrary. The dataset is created by first generating a 2D periodic Gaussian random field on a square domain $0,1~\text{mm}]^{2}$. The Marching Squares algorithm is then applied to extract the contour at a random threshold value, defining the boundary of the 2D periodic unit cell, with field values below the threshold treated as void.
@@
The predicted solutions at the final strain step are shown in Fig.˜6, comparing the ground truth with GINOT predictions. The first row highlights the median case with an $L_{2}$ relative error of 7.54%, while the second row shows the 95th percentile case with an $L_{2}$ relative error of 13.03%. The first column illustrates the input surface point cloud, the second column presents the true stress on the true deformed shape, the third column displays the predicted stress on the predicted deformed shape, and the last column shows the absolute error of stress on the true deformed shape. The mean $L_{2}$ relative error across the testing samples is 9.05%, with a standard deviation of 3.53%.
The stress and displacement solutions over historical strain steps are compared with the ground truth and illustrated in Fig.˜7. For brevity, only the median case is presented here. Additional cases, such as the best and worst, can be found in our Github repository. The first row shows the true stress on the true deformed shape, the second row depicts the predicted stress on the predicted deformed shape, and the last row highlights the absolute error of stress on the true deformed shape. The overall $L_{2}$ relative error for historical steps is 9.08%. The mean $L_{2}$ relative error for the testing samples is 9.06%, with a standard deviation of 3.02%.
As detailed in Section˜2.1, the geometry encoder in GINOT is designed to handle 1) nonuniform point cloud densities, 2) permutations in point order, and 3) padding within the point cloud. The first property is evident throughout the examples in this work. To demonstrate the latter two properties, we analyze the second-best test case in this example. Notably, GINOT maintains stable inference performance even when the point cloud is made sparser. In this analysis, the point cloud is shuffled, reduced to 80% of its original size, padded with 199 points at (1000,1000)(-1000,-1000), and shuffled again. These modified point clouds are input to the trained GINOT, with results shown in Fig.˜8. The $L_{2}$ relative errors for the original, reduced and shuffled, padded, and shuffled padded point clouds are 3.75%, 3.83%, 3.75%, and 3.75%, respectively. The similar $L_{2}$ errors across these cases demonstrate the geometry encoder’s robustness to point order and padding. We emphasize that the results in Fig.˜8 were produced using sampling and grouping layers implemented to be invariant to point permutations as described in Section˜2.1; consequently, the padded and shuffled-padded variants yield identical $L_{2}$ errors (3.75%). If a simpler FPS implementation is used that initializes from the first non-padding point (default of GINOT), the shuffled-padded case shows an increase in $L_{2}$ of only 0.002 relative to the padded case, indicating negligible impact. Although shuffling and padding have minimal impact on model performance, a significant reduction in point cloud density can lead to an increase in the $L_{2}$ relative error due to the loss of geometric information. The effect of point cloud density on model performance is summarized in Table˜5.
Table 5: Impact of point cloud density on the inference performance of the trained model.
Point cloud density
(percentage of original)
100% 80% 60% 40% 20%
$L_{2}$ relative error 3.75% 3.83% 4.50% 8.18% 30.46%
@@ 3.6 Jet engine bracket
The final example demonstrating GINOT’s capability is the Jet Engine Bracket (JEB) dataset [hong2024deepjeb]. This dataset comprises 2,138 samples of jet engine brackets derived from the "GE Jet Engine Bracket Challenge" guidelines. These geometries are complex 3D structures with arbitrary shapes and sizes. FEM simulations provide solution fields (e.g., stress, displacement) under four loading scenarios: vertical, horizontal, diagonal, and torsional, all with fixed loading magnitudes. In this example, we focus on the Mises stress solution under vertical loading.
@@
Figure 11: Mises stress solutions for the testing samples of the JEB dataset. The first row shows the best case ($L_{2}=16.99\%$), the second row represents the median case ($L_{2}=29.0\%$), and the third row depicts the worst case ($L_{2}=54.86\%$). The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
@@
The dataset includes CAD models, volume meshes in VTK format, and solution fields on fine quadratic tetrahedral elements. Instead of extracting surface nodes from the volume mesh, we extract the boundary point cloud directly from the CAD model vertices for each JEB sample using FreeCAD with PythonAPI, as illustrated in Fig.˜9(a). To reduce computational costs, the volume mesh is downgraded from 10-node tetrahedrons to 4-node tetrahedrons by removing the middle nodes of each element’s edge. The nodes of these linear tetrahedrons are then used as query points in the solution decoder of GINOT, as shown in Fig.˜9(a). Fig.˜9(b) illustrates the number of boundary points in the cloud and the volume mesh nodes for the 2,138 JEB samples. The boundary point cloud contains between 2,253 and 10,388 points, while the volume mesh nodes range from 18,085 to 55,009.
@@
Fig.˜10 illustrates the training process of GINOT on the JEB dataset. The mean squared error (MSE) loss for both the training and validation datasets decreases steadily during training. However, the testing loss remains significantly higher than the training loss, suggesting overfitting. This behavior is attributed to GINOT’s high representational capacity—its geometry-aware encoder and transformer decoder can closely fit intricate training patterns. The JEB dataset is comparatively small (2,138 samples) yet contains highly complex geometries. This combination of high model capacity and limited, heterogeneous data increases susceptibility to overfitting.
@@
After training, we evaluate GINOT’s performance on both the training and testing samples. For the training samples, the mean $L_{2}$ relative error is 18.77% with a standard deviation of 4.13%, while for the testing samples, the mean $L_{2}$ relative error is 30.77% with a standard deviation of 8.86%. Fig.˜11 illustrates the Mises stress solutions for the testing samples predicted by GINOT, showcasing the best, median, and worst cases in terms of $L_{2}$ relative error. The best case achieves $L_{2}=17.0\%$, while the worst case has $L_{2}=54.0\%$.
To assess efficiency relative to SDF-based neural operators, the signed distance function was computed on a $128^{3}$ grid for first 3D test geometry (39k surface mesh triangles). The SDF evaluation required about $30\,\text{s}$ on an A100 GPU, whereas GINOT’s sampling and grouping step took about $9\,\text{ms}$. This speed advantage shows that the proposed sampling and grouping strategy is substantially more efficient than explicit SDF construction for complex 3D geometries.
3.7 Ablation study
Finally, we conducted an ablation study to evaluate the impact of different configurations of our model.
The geometry encoder in our model includes cross-attention blocks followed by self-attention blocks. To assess the importance of these components, we performed experiments on the 2D PUC dataset and the 3D lug dataset. The results, summarized in Table˜6, are presented in terms of $L_{2}$ relative errors.
The study reveals that the cross-attention block is crucial, as its removal significantly increases the $L_{2}$ error. This is because the cross-attention block effectively encodes global geometric features from the boundary point cloud. On the other hand, the self-attention block’s necessity varies: its removal increases the $L_{2}$ error for the PUC dataset but decreases it for the lug dataset. This discrepancy is likely attributed to the higher complexity and randomness of the 2D PUC dataset compared to the parametric 3D lug dataset, where the self-attention block aids in learning better geometric representations for the more complex dataset.
Table 6: Ablation study results showing the $L_{2}$ relative errors for different attention block configurations.
2D PUC $\%$] 3D LUG $\%$]
Full model 7.81 0.40
Without cross-attention block 12.50 0.42
Without self-attention block 10.77 0.384
Without both cross- and self-blocks 21.84 0.389
Table 7: Ablation study results showing the $L_{2}$ relative errors (%) for different sampling and grouping configurations: number of sampled centroids $N_{s}$, points per group $N_{p}$, and grouping radius $r$. "–" indicates that the configuration was not tested.
$N_{s}$ PUC Elasticity $N_{p}$ PUC Elasticity $r$ PUC Elasticity LUG JEB
8 1.09 8 7.81 1.27 0.1 7.61 1.25 0.41 30.77
16 8.99 1.03 18 8.33 1.33 0.2 7.81 1.33 0.43 30.84
32 8.27 1.16 32 7.46 1.36 0.4 8.98 1.30
64 7.97 1.33 64 7.40 1.25 0.5 9.23 1.53 0.40 31.70
96 1.49
128 7.81
Table 8: Ablation on PC density for dataset of PUC. Reported are $L_{2}$ relative errors (%) when evaluating with the same density as training or with the full (100%) point cloud.
Training PC density
(% of original)
$L_{2}$ when eval.
with same density
$L_{2}$ when eval.
with 100% PC
100% 7.94 7.94
80% 8.49 8.99
60% 9.65 9.44
40% 11.46 10.22
We further examine the impact of the sampling and grouping hyperparameters in Table 7. We vary the number of sampled centroids ($N_{s}$), the points per group ($N_{p}$), the grouping radius ($r$), and report the relative $L_{2}$ error on the PUC, Elasticity, LUG, and JEB datasets. For Elasticity (105 boundary points), changing percentage of sampling points ($N_{s}/105$) from 7.6% to 30% yields similar accuracy, while increasing it to 61-91% degrades performance, likely because infomation become overly global. For PUC (up to 484 boundary points), increasing $N_{s}/484$ from 3.3% to 26.4% steadily improves accuracy, with diminishing returns beyond roughly 20-26%. A practical choice is $N_{s}$ at about 15-30% of the boundary points. The effect of $N_{p}$ is relatively minor on both PUC and Elasticity within the tested range. The grouping radius $r$ controls neighborhood size: smaller radii (r=0.1r=0.1-$0.2$) generally perform best, whereas larger radii (r=0.4r=0.4-$0.5$) tend to blend unrelated regions and may slightly hurt accuracy. We therefore recommend $r\in[0.1,0.2]$ together with a moderate $N_{s}$ ratio.
We then studied the effect of point-cloud (PC) density. Performance varies nonlinearly with density. As shown in Table˜8, reducing the PC density from 100% to 80% leaves the relative $L_{2}$ error essentially unchanged, indicating robustness to moderate downsampling. When the density drops to 40%, errors increase markedly due to the loss of fine geometric detail needed to capture local structure. Thus, modest downsampling can reduce compute with minimal accuracy loss, but aggressive sparsification should be avoided. Evaluating with denser PCs than those used in training only slightly affects accuracy when the training PC density is high (e.g., $\geq 60\%$), but yields accuracy improvement when the training PC density is low (e.g., 40%). These results confirm that reducing point-cloud density, which removes geometric detail, degrades accuracy.